# Synthetic Data Generation & Multi-Step RL for Reasoning & Tool Use

### v1

@@ 2.1 Multi-Step Data Collection
In Stage 1 (see Figure 1), we generate synthetic trajectories consisting of multiple steps of reasoning and tool use, which we use as training data for the step-wise RL methodology described in the next section. To compile a large-scale collection of synthetic trajectories, we augment a language model with a tool (e.g., a search engine or calculator), and iteratively prompt the model to generate multi-step trajectories. At each step, the model is asked to choose whether to call a tool or produce a final answer, and is always free to generate chains of thought (which it typically does). If the model generates a tool use call, it is parsed from the overall response, executed in the environment, and the result is presented to the model in the next step. See Appendix A for the prompt, which contains a question, explicit instructions regarding multi-step tool utilization, and the results of prior tool use calls.
@@
Having compiled these datasets, we consider four different filtering strategies and measure their impact on performance (Figure 1): (1) No filtering; (2) Process filtering, where we retain trajectories in which each step was deemed reasonable given all previous steps. Concretely, a model (Gemini 1.5 Pro Thinking, in our case) is prompted to render a binary judgment as to whether action $a_{i}$ is reasonable given the context $s_{i}$. See Appendix A for our prompt. No golden labels are used; (3) Outcome filtering, where we select trajectories based solely on whether the final response, $a_{K}$, matches the golden answer; and (4) Process and outcome filtering, in which we take the intersection of both filtering approaches and retain only trajectories that exhibit both step-wise soundness and correct final outcomes.
@@ 2.2 Step-Wise Reinforcement Learning Methodology
Thus, our objective function is the expected sum of stepwise rewards:
@@ 2.3 Step-Wise Inference-time Evaluation
As shown in Figure 3, at inference time, we iteratively prompt the model to either call a tool or produce a final answer. If the model generates a search query (indicated by $<$search_query$>$ $<$/search_query$>$ tags), we parse out that query, embed it with a Gecko model, perform a nearest neighbor lookup in the corresponding vector database, and inject the retrieved article into the model’s context window. If the model generates a calculator tool call (indicated by $<$math_exp$>$ $<$/math_exp$>$ tags), we parse out the mathematical expression, execute it with a SymPy interpreter, and inject the calculated results into the context window. This process terminates when the model either produces an answer (signaled by producing $<$answer$>$ $<$/answer$>$ tags) or reaches the maximum number of queries (5 for question-answering datasets, and 10 for mathematical reasoning datasets). See Appendix D for example trajectories.
@@ 3 Related Work
Reinforcement Learning for LLM Finetuning. One prominent approach, Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Christiano et al., 2023), consists of training a reward model on human preference labels at the response level, followed by RL optimization using Proximal Policy Optimization (PPO) (Schulman et al., 2017). Building upon this framework, Reinforcement Learning with AI Feedback (RLAIF) (Bai et al., 2022) has emerged as a scalable alternative that leverages AI models to generate feedback based on predefined principles or constitutions, reducing the need for costly human annotations. RL from Execution Feedback (RLEF) (Gehring et al., 2025) uses environment feedback, such as pass rate on coding test cases, to calculate the reward, which it then optimized via PPO. Besides PPO, other RL optimizations, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) and its successors (e.g., Azar et al. (2023); Ethayarajh et al. (2024); Meng et al. (2024); Lanchantin et al. (2025)) as well as GRPO (Shao et al., 2024) have also proven to be effective for finetuning LLMs to maximize a target reward. A limitation of the above approaches is that they focus on single-step optimization with the reward being calculated only at the end of the episode, leading to suboptimal performance for multi-step optimization (Liu et al., 2024; Wang et al., 2024). In SWiRL, we focus on scenarios where multiple steps of reasoning and tool calls are necessary prior to generating a response. Unlike the above methods, SWiRL enables the model to receive feedback on its granular stepwise actions which leads to better multi-step reasoning and tool use across longer horizons.
@@ 4.1 Evaluation Datasets
For the experiments in Table 1, we follow the same procedure as Wu et al. (2024), evaluating performance on 300 randomly subsampled examples from the target dataset, using the same language model as a judge (GPT4o) and the same prompt. For every other experiment in this paper, we used Gemma-2-27b as our judge, as this was more cost effective, with the exception of GSM8K for which we used Gemini 1.5 Pro as it exhibited noticeably better numeric evaluation. Model-based evaluation is emerging as a scalable and less brittle alternative to exact match and F1 metrics (Zheng et al., 2023; Gu et al., 2025), but does introduce a new source of stochasticity into the evaluation. See Appendix C for our own manual inspection and error analysis of three different model judges.
@@ 4.2 Results and Discussion
Figure 4: Impact of Data Filtering on Model Performance. Synthetic data for training is derived from HotPotQA. SWiRL learns to perform multi-hop question answering even when trained on unfiltered synthetic data. SWiRL’s best performance comes from training on process-only filtered data, where the data is selected based on the soundness of each step within its reasoning traces, but which includes both correct and incorrect responses.
@@
Generalization Across Disparate Tasks: To measure generalization across training tasks, we evaluated the mathematical reasoning capabilities of a model trained on multi-hop question-answering with search tool use (HotPotQA). Specifically, we evaluated the performance of this model on GSM8K, a mathematical reasoning task, providing the model with a SymPy interpreter to use as a calculator. This experiment was run on a different random subsample of 300 examples. As shown in Table 2, applying SWiRL on out-of-distribution data and tasks still improves performance.
@@
Figure 5: Comparison of SFT and SWiRL. Synthetic data for training is derived from HotPotQA, and to derive accuracy, Gemma 2 27b evaluates whether the model’s answer matches the golden answer. SWiRL greatly benefits from process-only filtered traces, and unlike SFT, is capable of learning from traces with both correct and incorrect outcomes.
Comparison of Supervised Finetuning and SWiRL: Figure 5 compares the performance of Supervised Fine-Tuning (SFT) and SWiRL on downstream tasks. The results show that SFT leads to worse overall performance when compared to SWiRL across all data filtering strategies. We observe that SFT performs better if we apply it to data that is both process and outcome-filtered, rather only process-filtered. However, interestingly, SWiRL learns best from data that is only process-filtered. We attribute this to SFT’s tendency to memorize, rather than generalize (Chu et al., 2025; Setlur et al., 2024), which can hinder the model’s performance on new, unseen scenarios. In contrast, SWiRL has the ability to improve model performance by targeting per-step reward maximization. SWiRL enables the model to develop a deeper understanding of the necessary steps of query generation and retrieval, which leads to enhanced planning and generalization.
@@
Figure 6: Performance of SWiRL With and Without Multi-Step Tool Use. SWiRL’s multi-step tool use inference improves the performance of both the base model and the SWiRL-finetuned model, but benefits the latter substantially more. Even without access to a tool, the SWiRL model outperforms, suggesting that SWiRL training improves the model’s ability to break down complex problems into multiple manageable subtasks.
@@
Impact of Scaling Finetuning Dataset and Model Size: Our experiments on scaling the fine-tuning dataset size reveal a clear trend: SWiRL has the ability to leverage larger datasets, even when using only process-filtered data, as shown in Figure 7. As the fine-tuning dataset size increases, a consistent enhancement in model performance is observed across our target multi-step reasoning tasks. While a limited dataset of 100 data points appears insufficient for the model to effectively generalize, a significant improvement is evident with 1,000 data points, showing solid gains across all datasets. Furthermore, scaling up to 10,000 data points continues to yield further performance enhancements, confirming the efficacy of our method in capitalizing on larger datasets for improved reasoning capabilities.
@@
We also varied model size, observing that smaller models (2b and 9b) may benefit from in-domain SWiRL, but do not display the same generalization as their larger counterpart, Gemma 2 27b. See results in Appendix B.
Effect on Mean Process Label Accuracy:
@@
Figure 7: Performance as a Function of Synthetic Dataset Size. Synthetic training data is derived from HotPotQA, and accuracy is evaluated by Gemma 2 27b. As we scale the dataset size, we observe consistent improvements in model performance. With only 1000 data points, the model robustly improves both on in- and out- of distribution datasets.
@@
In the previous subsections, we evaluated the effect of SWiRL on downstream task accuracy. Here, we take a deeper look to understand how SWiRL achieves these performance improvements. In Table 3, we show the average process label accuracy for the baseline model vs. a SWiRL finetuned model on 500 trajectories (seeded by 100 questions) for both HotPotQA and GSM8K. To calculate the score per step, we use the same model and prompt as we used for process filtering, as described in Section 4. We take a macro-average of the process label scores within and then across trajectories. We observe that both for in-distribution and out-of-distribution tasks, the SWiRL model generates trajectories with higher average process labels, suggesting that the higher final accuracies are driven by better multi-step reasoning.
@@ 5 Conclusion
Appendix A Prompts for Synthetic Data Generation, Filtering, and Evaluation
@@
Appendix B Impact of Model Size on Effectiveness of SWiRL.
Appendix C Error Analysis of Three LLM Judges
Table 4: Error Rates for Gemma-2-27b Judgments on HotPotQA (N=100)
Metric Rate (%)
False Positive Rate (FPR) $4$
False Negative Rate (FNR) $1$
Table 5: Manual Analysis of LLM Math Grading Accuracy (N=100)
Model FPR FNR Notes
GPT-4o 0 10 Overly harsh; all errors involved units.
Appendix D Example Trajectories

### v2

@@ 2.1 Multi-Step Data Collection
In Stage 1 (see Figure 1), we generate synthetic trajectories consisting of multiple steps of reasoning and tool use, which we use as training data for the step-wise RL methodology described in the next section. To compile a large-scale collection of synthetic trajectories, we augment a language model with a tool (e.g., a search engine or calculator), and iteratively prompt the model to generate multi-step trajectories. At each step, the model is asked to choose whether to call a tool or produce a final answer, and is always free to generate chains of thought (which it typically does). If the model generates a tool use call, it is parsed from the overall response, executed in the environment, and the result is presented to the model in the next step. See Appendix E for the prompt, which contains a question, explicit instructions regarding multi-step tool utilization, and the results of prior tool use calls.
@@
Having compiled these datasets, we consider four different filtering strategies and measure their impact on performance (Figure 1): (1) No filtering; (2) Process filtering, where we retain trajectories in which each step was deemed reasonable given all previous steps. Concretely, a model (Gemini 1.5 Pro Thinking, in our case) is prompted to render a binary judgment as to whether action $a_{i}$ is reasonable given the context $s_{i}$. See Appendix E for our prompt. No golden labels are used; (3) Outcome filtering, where we select trajectories based solely on whether the final response, $a_{K}$, matches the golden answer; and (4) Process and outcome filtering, in which we take the intersection of both filtering approaches and retain only trajectories that exhibit both step-wise soundness and correct final outcomes.
@@ 2.2 Step-Wise Reinforcement Learning Methodology
Thus, our objective function is the expected sum of step-wise rewards:
@@ 2.3 Step-Wise Inference-time Evaluation
As shown in Figure 3, at inference time, we iteratively prompt the model to either call a tool or produce a final answer. If the model generates a search query (indicated by $<$search_query$>$ $<$/search_query$>$ tags), we parse out that query, embed it with a Gecko model, perform a nearest neighbor lookup in the corresponding vector database, and inject the retrieved article into the model’s context window. If the model generates a calculator tool call (indicated by $<$math_exp$>$ $<$/math_exp$>$ tags), we parse out the mathematical expression, execute it with a SymPy interpreter, and inject the calculated results into the context window. This process terminates when the model either produces an answer (signaled by producing $<$answer$>$ $<$/answer$>$ tags) or reaches the maximum number of queries (5 for question-answering datasets, and 10 for mathematical reasoning datasets). See Appendix F for example trajectories.
@@ 3 Related Work
Reinforcement Learning for LLM Finetuning. One prominent approach, Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Christiano et al., 2023), consists of training a reward model on human preference labels at the response level, followed by RL optimization using Proximal Policy Optimization (PPO) (Schulman et al., 2017). Building upon this framework, Reinforcement Learning with AI Feedback (RLAIF) (Bai et al., 2022) has emerged as a scalable alternative that leverages AI models to generate feedback based on predefined principles or constitutions, reducing the need for costly human annotations. RL from Execution Feedback (RLEF) (Gehring et al., 2025) uses environment feedback, such as pass rate on coding test cases, to calculate the reward, which it then optimized via PPO. Besides PPO, other RL optimizations, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023) and its successors (e.g., Azar et al. (2023); Ethayarajh et al. (2024); Meng et al. (2024); Lanchantin et al. (2025)) as well as GRPO (Shao et al., 2024) have also proven to be effective for finetuning LLMs to maximize a target reward. A limitation of the above approaches is that they focus on single-step optimization with the reward being calculated only at the end of the episode, leading to suboptimal performance for multi-step optimization (Liu et al., 2024; Wang et al., 2024). In SWiRL, we focus on scenarios where multiple steps of reasoning and tool calls are necessary prior to generating a response. Unlike the above methods, SWiRL enables the model to receive feedback on its granular step-wise actions which leads to better multi-step reasoning and tool use across longer horizons.
@@ 4.1 Evaluation Datasets
For the experiments in Table 1, we follow the same procedure as Wu et al. (2024), evaluating performance on 300 randomly subsampled examples from the target dataset, using the same language model as a judge (GPT4o) and the same prompt. For every other experiment in this paper, we used Gemma-2-27b as our judge, as this was more cost effective, with the exception of GSM8K for which we used Gemini 1.5 Pro as it exhibited noticeably better numeric evaluation. Model-based evaluation is emerging as a scalable and less brittle alternative to exact match and F1 metrics (Zheng et al., 2023; Gu et al., 2025), but does introduce a new source of stochasticity into the evaluation. See Appendix D for our own manual inspection and error analysis of three different model judges.
@@ 4.2 Results and Discussion
Figure 4: Impact of Data Filtering on SWiRL’s Performance. Synthetic data for training is derived from HotPotQA. SWiRL learns to perform multi-hop question answering even when trained on unfiltered synthetic data. SWiRL’s best performance comes from training on process-only filtered data, where the data is selected based on the soundness of each step within its reasoning traces, but which includes both correct and incorrect responses. Note that in all cases, the model is provided with the tool and is allowed to call it multiple times.
@@
Generalization Across Disparate Tasks and Tools: To measure generalization across training tasks, we evaluated the mathematical reasoning capabilities of a model trained on multi-hop question-answering with search tool use (HotPotQA). Specifically, we evaluated the performance of this model on GSM8K, a mathematical reasoning task, providing the model with a SymPy interpreter to use as a calculator. This experiment was run on a different random subsample of 300 examples. As shown in Table 2, applying SWiRL on out-of-distribution data and tasks still improves performance.
@@
Figure 5: Comparison of SFT and SWiRL. SWiRL greatly benefits from process-only filtered traces, and unlike SFT, is capable of learning from traces with both correct and incorrect outcomes. Note that in all cases, the models have access to the same tool (retriever or calculator) and are allowed to call it multiple times.
Comparison of Supervised Finetuning and SWiRL: Figure 5 compares the performance of Supervised Fine-Tuning (SFT) and SWiRL on various downstream tasks. The results show that SFT leads to worse overall performance compared to SWiRL. In fact, we found that SFT actually degrades performance compared to the base model (see Appendix B), which is consistent with prior work showing that SFT can harm reasoning capabilities (Chen et al., 2025). Interestingly, we observe that SFT performs better if we apply it to data that is both process and outcome-filtered, rather only process-filtered, whereas SWiRL learns best from data that is only process-filtered. We attribute this to SFT’s tendency to memorize, rather than generalize (Chu et al., 2025; Setlur et al., 2024), which can hinder the model’s performance on new, unseen scenarios. In contrast, SWiRL has the ability to improve model performance by targeting per-step reward maximization. SWiRL enables the model to develop a deeper understanding of the necessary steps (e.g. multiple steps of query generation and retrieval), which leads to enhanced planning and generalization. Additionally, in Appendix B, we ran SFT on synthetic trajectories generated by Gemini 1.5 Pro (the reward model in SWiRL), but found that this did not improve performance.
@@
Figure 6: Performance of SWiRL With and Without Multi-Step Tool Use. SWiRL’s multi-step tool use inference improves the performance of both the base model and the SWiRL-finetuned model, but benefits the latter substantially more.
@@
Impact of Scaling Finetuning Dataset and Model Size: Our experiments on scaling the fine-tuning dataset size reveal a clear trend: SWiRL has the ability to leverage larger datasets, even when using only process-filtered data, as shown in Figure 7. As the fine-tuning dataset size increases, a consistent enhancement in model performance is observed across our target multi-step reasoning tasks. While a limited dataset of 100 data points appears insufficient for the model to effectively generalize, a significant improvement is evident with 1,000 data points, showing solid gains across all datasets. Furthermore, scaling up to 10,000 data points continues to yield further performance enhancements, confirming the efficacy of our method in capitalizing on larger datasets for improved reasoning capabilities. Interestingly, in GSM8k, we also observe that performance improves as we scale up the number of synthetic trajectories, even though these trajectories represent a different domain (question-answering vs. mathematical reasoning) and a different tool (search vs. calculator).
@@
We also varied model size, observing that smaller models (2b and 9b) may benefit from in-domain SWiRL, but do not display the same generalization as their larger counterpart, Gemma-2-27b. See results in Appendix C.
@@
Figure 7: Performance as a Function of Synthetic Dataset Size. Synthetic training data is derived from HotPotQA, and accuracy is evaluated by Gemma-2-27b. As we scale the dataset size, we observe consistent improvements in model performance. With only 1000 data points, the model robustly improves both on in- and out- of distribution datasets. In all cases, multi-step SWiRL inference during evaluation.
Comparison Against the Reward Model: Here, we compare the performance of Gemma-2-27b before and after being finetuned via SWiRL against the reward model used in SWiRL, Gemini 1.5 Pro. Similar to other experiments in this paper, Gemini-2-27b is used for synthetic data generation and Gemini 1.5 Pro is used as the reward model. The results are shown in Figure 8. As can be seen, SWiRL significantly outperforms the base model on all benchmarks and even outperforms Gemini 1.5 Pro on some out-of-distribution benchmarks, including CofCA and BeerQA. These results suggest that SWiRL is not merely distilling a stronger reward model (Gemini 1.5 Pro).
Figure 8: Comparing the performance of SWiRL with the base model and Gemini 1.5 Pro. Finetuning with SWiRL improves performance on all benchmarks, and even enables the model to outperform Gemini 1.5 Pro on some out-of-distribution benchmarks, suggesting that SWiRL is not merely distilling the larger reward model (Gemini 1.5 Pro). Note that, in all cases, each model has access to the same tool (retriever or calculator) and is allowed to call it multiple times.
@@
Effect on Mean Process Label Accuracy: In previous subsections, we evaluated the effect of SWiRL on downstream task accuracy. Here, we examine how SWiRL achieves these performance improvements. In Table 3, we show the average process label accuracy for the baseline model vs. a SWiRL finetuned model on 500 trajectories (seeded by 100 questions) for both HotPotQA and GSM8K. To calculate the score per step, we use the same model and prompt as we used for process filtering, as described in Section 4. We take a macro-average of the process label scores within and then across trajectories. We observe that both for in-distribution and out-of-distribution tasks, the SWiRL model generates trajectories with higher average process labels, suggesting that the higher final accuracies are driven by better multi-step reasoning.
@@ 5 Conclusion
Appendix A Additional Evaluation
Statistical Analysis of SWiRL’s Performance Improvements: To ensure that the improvements from SWiRL were meaningful, we evaluated on a larger number of samples, and reported traditional metrics like F1, Precision, and Recall. As shown in the table below, SWiRL improves performance across all benchmarks and all metrics, and by a margin that is substantially larger than the margin of error (95% confidence interval).
Dataset Model Model Score F1 Score Precision Recall
BeerQA Base $0.597\,\text{$\,\pm\,$0.025}$ $0.470\,\text{$\,\pm\,$0.025}$ $0.458\,\text{$\,\pm\,$0.025}$ $0.565\,\text{$\,\pm\,$0.025}$
SWiRL $0.667\,\text{$\,\pm\,$0.024}$ $0.561\,\text{$\,\pm\,$0.025}$ $0.557\,\text{$\,\pm\,$0.025}$ $0.613\,\text{$\,\pm\,$0.025}$
CofCA Base $0.583\,\text{$\,\pm\,$0.032}$ $0.363\,\text{$\,\pm\,$0.031}$ $0.342\,\text{$\,\pm\,$0.031}$ $0.552\,\text{$\,\pm\,$0.033}$
SWiRL $0.648\,\text{$\,\pm\,$0.031}$ $0.467\,\text{$\,\pm\,$0.033}$ $0.468\,\text{$\,\pm\,$0.033}$ $0.573\,\text{$\,\pm\,$0.032}$
HotpotQA Base $0.667\,\text{$\,\pm\,$0.024}$ $0.577\,\text{$\,\pm\,$0.025}$ $0.577\,\text{$\,\pm\,$0.025}$ $0.633\,\text{$\,\pm\,$0.024}$
SWiRL $0.747\,\text{$\,\pm\,$0.022}$ $0.678\,\text{$\,\pm\,$0.024}$ $0.684\,\text{$\,\pm\,$0.024}$ $0.705\,\text{$\,\pm\,$0.023}$
MuSiQue Base $0.431\,\text{$\,\pm\,$0.025}$ $0.305\,\text{$\,\pm\,$0.023}$ $0.304\,\text{$\,\pm\,$0.023}$ $0.364\,\text{$\,\pm\,$0.024}$
SWiRL $0.509\,\text{$\,\pm\,$0.025}$ $0.398\,\text{$\,\pm\,$0.025}$ $0.397\,\text{$\,\pm\,$0.025}$ $0.434\,\text{$\,\pm\,$0.025}$
GSM8k Base $0.685\,\text{$\,\pm\,$0.024}$ $0.337\,\text{$\,\pm\,$0.024}$ $0.311\,\text{$\,\pm\,$0.023}$ $0.445\,\text{$\,\pm\,$0.025}$
SWiRL $0.765\,\text{$\,\pm\,$0.021}$ $0.500\,\text{$\,\pm\,$0.025}$ $0.480\,\text{$\,\pm\,$0.025}$ $0.550\,\text{$\,\pm\,$0.025}$
Table 4: Performance Comparison of Base vs. SWiRL Models. Values formatted as Mean $\pm$ Margin of Error for a 95% confidence interval. Sample size $n=1500$ for all datasets except for CofCA for which we reported results on all examples for $n=900$. To calculate margin of error, we estimate standard deviation as $\sqrt{p(1-p)}$ where $p$ is the mean score.
Appendix B Impact of Synthetic Data Source on SFT Performance
Figure 9: Impact of Synthetic Data Source on SFT Performance. In an attempt to level the playing field between SFT and SWiRL, we performed SFT with trajectories generated by Gemini 1.5 Pro (the reward model in SWiRL), but this did not improve its performance. The training data consists of synthetic trajectories seeded by HotPotQA questions.
We found that performing SFT on multi-step synthetic reasoning traces actually degraded performance, an observation which is consistent with other recent work (Chen et al., 2025; Chu et al., 2025; Setlur et al., 2024). However, since we used a stronger model (Gemini 1.5 Pro) as a reward model in SWiRL (which SFT did not have access to), we investigated whether we could improve SFT performance by using trajectories from Gemini 1.5 Pro. Note that it is more costly to generate trajectories from Gemini 1.5 Pro compared to merely using it as a reward signal. We observed that Gemma’s performance further degrades if we perform supervised finetuning on trajectories from Gemini instead of Gemma. This could be due to the fact that Gemini has a different data distribution from Gemma, causing it to be more difficult for Gemma to learn from these out-of-distribution reasoning traces.
Appendix C Impact of Model Size on Effectiveness of SWiRL.
Appendix D Error Analysis of Three LLM Judges
Table 5: Error Rates for Gemma-2-27b Judgments on HotPotQA (N=100)
Metric Rate (%)
False Positive Rate (FPR) $4$
False Negative Rate (FNR) $1$
Table 6: Manual Analysis of LLM Math Grading Accuracy (N=100)
Model FPR FNR Notes
GPT-4o 0 10 Overly harsh; all errors involved units.
Appendix E Prompts for Synthetic Data Generation, Filtering, and Evaluation
@@
Appendix F Example Trajectories